# Off-policy Evaluation with Deeply-abstracted States

### v2

@@ 3.1 Data generating process, policy, value and IS ratio
Here, the numerator represents the discounted probability of visiting a given state-action pair under the target policy $\pi$, a crucial component in policy-based learning for estimating $\pi^{*}$ (Sutton et al.,, 1999; Schulman et al.,, 2015). The denominator corresponds to the limiting state-action distribution under the behavior policy. These ratios are fundamental in constructing IS estimators, as detailed in Methods 2 and 3 of Section 3.3.
3.2 State abstractions for policy learning
Let $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho_{0},\gamma\rangle$ be the ground MDP. A state abstraction $\phi$ is a mapping from the state space $\mathcal{S}$ to certain abstract state space $\mathcal{X}=\{\phi(s):s\in\mathcal{S}\}$. Below, we review some commonly studied definitions of state abstraction designed for learning the optimal policy $\pi^{*}$; see Jiang, (2018).
@@ 3.2 State abstractions for policy learning
Definition 2 ($Q^{*}$-irrelevance)
$\phi$ is $Q^{*}$-irrelevant if for any $s^{(1)}$, $s^{(2)}\in\mathcal{S}$ whenever $\phi(s^{(1)})=\phi(s^{(2)})$, the optimal Q-function satisfies $Q^{*}(a,s^{(1)})=Q^{*}(a,s^{(2)})$ for any $a\in\mathcal{A}$.
Definitions 1 and 2 are easy to understand, requiring the optimal policy/Q-function to depend on a state $s$ only through its abstraction $\phi(s)$. In practical terms, these definitions encourage the transformation of raw MDP data into a new sequence of state-action-reward triplets $(\phi(S),A,R)$ for policy learning. However, the transformed data may not necessarily satisfy the Markov assumption. This leads us to define the following model-irrelevance, which aims to preserve the MDP structure while ensuring $\pi^{*}$- and $Q^{*}$-irrelevance.
@@
The first condition in (3) corresponds to “reward-irrelevance” whereas the second condition represents “transition-irrelevance”. Consequently, Definition 3 defines a “model-based” abstraction, in contrast to “model-free” abstractions considered in Definitions 1 and 2. Notice that the term $\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathcal{T}(s^{\prime}|a,s)$ – appearing in the second equation of (3) – represents the probability of transitioning to $\phi(S^{\prime})=x^{\prime}$ in the abstract state space. Thus, the second condition essentially requires the abstract next state $\phi(S^{\prime})$ to be conditionally independent of $S$ given $A$ and $\phi(S)$. Assuming $S$ can be decomposed into the union of $\phi(S)$ and $\psi(S)$, which represent relevant features and irrelevant features, respectively. This condition implies that the evolution of those relevant features depends solely on themselves, independent of those irrelevant features. This ensures that the transformed data triplets $(\phi(S),A,R)$ remains an MDP. Meanwhile, the evolution of those irrelevant features may still depend on the relevant features; see Figure 1(a) for an illustration.
It is also known that model-irrelevance implies $Q^{*}$-irrelevance, which in turn implies $\pi^{*}$-irrelevance; see e.g., Theorem 2 in Li et al., (2006). Given that the transformed data remains an MDP under model-irrelevance, one can apply existing state-of-the-art RL algorithms to the abstract state space instead of the original ground space, leading to more effective learning of the optimal policy.
Figure 1: Illustrations of (a) model-irrelevance and (b) backward-model-irrelevance. $S_{t}$ is decomposed into the union of $\phi(S_{t})$ (relevant features) and $\psi(S_{t})$ (irrelevant features).
3.3 OPE methodologies
We focus on four OPE methods, covering the three families of estimators introduced in Section 2. Each method employs a specific formula to identify $J(\pi)$, which we detail below. The first method is a popular value-based approach – the Q-function-based method. The second and third methods are the two major IS estimators: SIS and MIS. The fourth method is a semi-parametrically efficient doubly robust method, double RL (DRL), known for achieving the smallest possible MSE among a broad class of OPE estimators (Kallus and Uehara,, 2020, 2022; Liao et al.,, 2022).
Method 1 (Q-function-based method). For a given Q-function $Q$, define $f_{1}(Q)$ as the estimating function $\sum_{a\in\mathcal{A}}\pi(a|S_{1})Q(a,S_{1})$ with $S_{1}$ being the initial state. By (1) and the definition of $J(\pi)$, it is immediate to see that $J(\pi)=\mathbb{E}[f_{1}(Q^{\pi})]$. This motivates the Q-function-based method which uses a plug-in estimator to approximate $\mathbb{E}[f_{1}(Q^{\pi})]$ and estimate $J(\pi)$. In particular, $Q^{\pi}$ can be estimated by Q-learning type algorithms (e.g., fitted Q-evaluation, FQE, Le et al.,, 2019), and the expectation can be approximated based on the empirical initial state distribution.
Method 2 (Sequential importance sampling). For a given IS ratio $\rho^{\pi}$, let $\rho_{1:t}^{\pi}$ denote the sequential IS ratio $\prod_{j=1}^{t}\rho^{\pi}(A_{j},S_{j})$. It follows from the change of measure theorem that the counterfactual reward $\mathbb{E}^{\pi}(R_{t})$ is equivalent to $\mathbb{E}(\rho^{\pi}_{1:t}R_{t})$ whose expectation is taken with respect to the offline data distribution. Assuming all trajectories in $\mathcal{D}$ terminate after a finite time $T$, this allows us to represent $J(\pi)$ by $\mathbb{E}[f_{2}(\rho^{\pi})]$ where $f_{2}(\rho^{\pi})=\sum_{t=1}^{T}\gamma^{t-1}\rho_{1:t}^{\pi}R_{t}$. SIS utilizes a plug-in estimator to first estimate $\rho^{\pi}$ (when the behavior policy is unknown), and then employs this estimator, along with the empirical data distribution, to approximate $\mathbb{E}[f_{2}(\rho^{\pi})]$. However, a notable limitation of this estimator is its rapidly increasing variance due to the use of the SIS ratio $\rho^{\pi}_{1:t}$. Specifically, this variance tends to grow exponentially with respect to $t$, a phenomenon often referred to as the curse of horizon (Liu et al.,, 2018).
Method 3 (Marginalized importance sampling). The MIS estimator is designed to overcome the limitations of the SIS estimator. It breaks the curse of horizon by taking the structure of the MDP model into account. As noted previously, under the Markov assumption, the reward depends only on the current state-action pair, rather than the entire history. This insight allows us to replace the SIS ratio with the MIS ratio, which depends solely on the current state-action pair. This modification considerably reduces variance because $w^{\pi}$ is no longer history-dependent. Assuming the data trajectory is stationary over time – that is, all state-action-reward $(S,A,R)$ triplets have the same distribution – it can be shown that $J(\pi)=\mathbb{E}[f_{3}(w^{\pi})]$ where $f_{3}(w^{\pi})=(1-\gamma)^{-1}w^{\pi}(A,S)R$ for any triplet $(S,A,R)$. Both $w^{\pi}$ and the expectation can be effectively approximated using offline data.
Method 4 (Double reinforcement learning). DRL combines the Q-function-based method with MIS. Let $f_{4}(Q,w)=f_{1}(Q)+(1-\gamma)^{-1}w(A,S)[R+\gamma\sum_{a}\pi(a|S^{\prime})Q(a,S^{\prime})-Q(A,S)]$, where $f_{1}$ is defined in Method 1 and $(S,A,R,S^{\prime})$ denotes a state-action-reward-next-state tuple. Under the stationarity assumption, it can be shown that $J(\pi)=\mathbb{E}[f_{4}(Q,w)]$ when either $Q=Q^{\pi}$ or $w=w^{\pi}$ (Kallus and Uehara,, 2022). DRL proposes to learn both $Q^{\pi}$ and $w^{\pi}$ from the data, employing these estimators to calculate $\mathbb{E}[f_{4}(Q,w)]$ and approximate the expectation with empirical data distribution. The resulting estimator benefits from double robustness: it is consistent when either $Q^{\pi}$ or $w^{\pi}$ is correctly specified.
4 Proposed state abstractions for policy evaluation
This section presents model-free (Section 4.1), model-based irrelevance conditions (Section 4.2) for OPE and analyzes the OPE estimators under these conditions (Lemma 1 & Theorem 1). Motivated by this analysis, we propose our iterative procedure (Section 4.3) and study its property (Theorem 2).
Our theoretical analysis is concerned with the Fisher consistency of various OPE estimators, named after, Ronald Fisher, the founder of modern statistics. In particular, the Fisher consistency requires an estimator to be exactly equal to the ground truth given infinite samples. Specialized to our settings, it imposes two requirements:
  1. (i)
    The identification formulas presented in Section 3.3 remain valid when replacing the oracle Q-function or (M)IS ratio with those projected into the proposed abstract state space;
  2. (ii)
    The Q-function or (M)IS ratio defined on the abstract state space is identifiable.
Once the Fisher consistency is established, the estimator’s finite sample properties can be readily obtained using existing techniques (see e.g., Uehara et al., 2021a, ). Therefore, for conciseness and to avoid redundancy, we chose not to present finite sample results in our theoretical analysis.
4.1 Model-free irrelevance conditions
We first introduce several model-free irrelevance conditions tailored for OPE.
Definition 4 ($\pi$-irrelevance)
Definition 5 ($Q^{\pi}$-irrelevance)
Definitions 4 and 5 are adaptations of Definitions 1 and 2 designed for policy evaluation, with the optimal policy $\pi^{*}$ replaced by the target policy $\pi$. The following definitions are tailored for IS estimators (see Methods 2 and 3 in Section 3.3).
Definition 6 ($\rho^{\pi}$-irrelevance)
Definition 7 ($w^{\pi}$-irrelevance)
These irrelevance conditions encourage us to conduct OPE on the abstract state space to reduce sample complexity. Nevertheless, methods for deriving abstractions that satisfy these conditions (particularly $Q^{\pi}$- and $w^{\pi}$-irrelevance) remain unclear. Furthermore, the state-action-reward triplets transformed via these abstractions $(\phi(S),A,R)$ might not maintain the MDP structure. This complicates the process of learning $Q_{\phi}^{\pi}$ and $w_{\phi}^{\pi}$. These challenges motivate us to consider model-based irrelevance conditions introduced in the subsequent section.
4.2 Model-based irrelevance conditions
@@
To begin with, we discuss two perspectives of the data generated within the MDP framework; see Figure 2 for a graphical illustration.
@@
    The first perspective is the traditional forward MDP model with all state-action-reward triplets sequenced by time index. This yields the model-based irrelevance condition defined in Definition 3. We will discuss the relationship between this condition and Definitions 5-7 below.
@@
    The second perspective offers a backward view by reversing the time order. Specifically, he symmetric nature of the Markov assumption — implying that if the future is independent of the past given the present, the past must also be independent of the future given the present — the reversed state-action pairs also maintain the Markov property. Leveraging this property, we define another backward MDP, which forms the basis for deriving model-based conditions for achieving $w^{\pi}$-irrelevance and motivates the subsequent iterative procedure. This development represents one of our main contributions.

### v3

@@ 3.1 Data generating process, policy, value and IS ratio
Here, the numerator represents the discounted probability of visiting a given state-action pair under the target policy $\pi$ – a crucial component in policy-based learning for estimating the optimal policy (denoted by $\pi^{*}$) (Sutton et al., 1999; Schulman et al., 2015). The denominator corresponds to the limiting state-action distribution under the behavior policy. These ratios are fundamental in constructing IS estimators, as detailed in Methods 2 and 3 of Section 3.2.
3.2 OPE methodologies
We focus on four OPE methods, covering the three families of estimators introduced in Section 2. Each method employs a specific formula to identify $J(\pi)$, which we detail below. The first method is a popular value-based approach – the Q-function-based method. The second and third methods are the two major IS estimators: SIS and MIS. The fourth method is a semi-parametrically efficient doubly robust method, double RL (DRL), known for achieving the smallest possible MSE among a broad class of OPE estimators (Kallus and Uehara, 2020, 2022; Liao et al., 2022).
Method 1 (Q-function-based method). For a given Q-function $Q$, define $f_{1}(Q)$ as the estimating function $\sum_{a\in\mathcal{A}}\pi(a|S_{1})Q(a,S_{1})$ with $S_{1}$ being the initial state. By (1) and the definition of $J(\pi)$, it is immediate to see that $J(\pi)=\mathbb{E}[f_{1}(Q^{\pi})]$. This motivates the Q-function-based method which uses a plug-in estimator to approximate $\mathbb{E}[f_{1}(Q^{\pi})]$ and estimate $J(\pi)$. In particular, $Q^{\pi}$ can be estimated by Q-learning type algorithms such as FQE, and the expectation can be approximated via the empirical initial state distribution.
Method 2 (Sequential importance sampling). For a given IS ratio $\rho^{\pi}$, let $\rho_{1:t}^{\pi}$ denote the sequential IS ratio $\prod_{j=1}^{t}\rho^{\pi}(A_{j},S_{j})$. It follows from the change of measure theorem that the counterfactual reward $\mathbb{E}^{\pi}(R_{t})$ is equivalent to $\mathbb{E}(\rho^{\pi}_{1:t}R_{t})$ whose expectation is taken with respect to the offline data distribution. Assuming all trajectories in $\mathcal{D}$ terminate after a finite time $T$, this allows us to represent $J(\pi)$ by $\mathbb{E}[f_{2}(\rho^{\pi})]$ where $f_{2}(\rho^{\pi})=\sum_{t=1}^{T}\gamma^{t-1}\rho_{1:t}^{\pi}R_{t}$. SIS utilizes a plug-in estimator to first estimate $\rho^{\pi}$ (when the behavior policy is unknown), and then employs this estimator, along with the empirical data distribution, to approximate $\mathbb{E}[f_{2}(\rho^{\pi})]$. However, a notable limitation of this estimator is its rapidly increasing variance due to the use of the SIS ratio $\rho^{\pi}_{1:t}$. Specifically, this variance tends to grow exponentially with respect to $t$, a phenomenon referred to as the curse of horizon (Liu et al., 2018).
Method 3 (Marginalized importance sampling). The MIS estimator is designed to overcome the limitations of the SIS estimator. It breaks the curse of horizon by taking the structure of the MDP model into account. As noted previously, under the Markov assumption, the reward depends only on the current state-action pair, rather than the entire history. This insight allows us to replace the SIS ratio with the MIS ratio, which depends solely on the current state-action pair. This modification considerably reduces variance because $w^{\pi}$ is no longer history-dependent. Assuming the data trajectory is stationary over time – that is, all state-action-reward $(S,A,R)$ triplets have the same distribution – it can be shown that $J(\pi)=\mathbb{E}[f_{3}(w^{\pi})]$ where $f_{3}(w^{\pi})=(1-\gamma)^{-1}w^{\pi}(A,S)R$ for any triplet $(S,A,R)$. Both $w^{\pi}$ and the expectation can be approximated using offline data.
Method 4 (Double reinforcement learning). DRL combines the Q-function-based method with MIS. Let $f_{4}(Q,w)=f_{1}(Q)+(1-\gamma)^{-1}w(A,S)[R+\gamma\sum_{a}\pi(a|S^{\prime})Q(a,S^{\prime})-Q(A,S)]$, where $f_{1}$ is defined in Method 1 and $(S,A,R,S^{\prime})$ denotes a state-action-reward-next-state tuple. Under the stationarity assumption, it can be shown that $J(\pi)=\mathbb{E}[f_{4}(Q,w)]$ when either $Q=Q^{\pi}$ or $w=w^{\pi}$ (Kallus and Uehara, 2022). DRL proposes to learn both $Q^{\pi}$ and $w^{\pi}$ from the data, employ these estimators to calculate $\mathbb{E}[f_{4}(Q,w)]$ and approximate the expectation with the empirical data distribution. The resulting estimator benefits from double robustness: it is consistent when either $Q^{\pi}$ or $w^{\pi}$ is correctly specified.
4 Proposed state abstractions for policy evaluation
This section presents our proposed iterative procedure for state abstraction. We begin by offering a bias-variance decomposition for the MSE of various post-abstraction OPE estimators, i.e., estimators applied to a given abstract space, and demonstrate that when the abstraction is an MSA, it generally reduces the MSE of the OPE estimator. Meanwhile, the finer the abstraction, the smaller the MSE (see Section 4.1). Inspired by this analysis, we put forward our iterative procedure (in Section 4.2).
4.1 A unified theory for post-abstraction OPE estimators
Recall that $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho_{0},\gamma\rangle$ denotes the ground MDP. An abstraction $\phi$ is a mapping from the state space $\mathcal{S}$ to an abstract state space $\mathcal{X}=\{\phi(s):s\in\mathcal{S}\}$. It transforms the raw MDP data into a new sequence of abstract state-action-reward-next-state tuples $(\phi(S),A,R,\phi(S^{\prime}))$. Recall that $\pi$ represents the target policy we aim to evaluate. Our analysis focuses on a specific subclass of abstractions that are $\pi$-irrelevant and preserve the Markov property, known as Markov state abstractions (Allen et al., 2021), and their formal definitions are provided below.
Definition 1 ($\pi$-irrelevance)
Definition 2 (Markov State Abstraction)
$\phi$ is an MSA if and only if both the transition and reward function on the transformed data satisfy the Markov property, i.e., for any $t$ and $x$, we have
$\mathbb{E}[R_{t}|A_{t},\phi(S_{t}),A_{t-1},\phi(S_{t-1}),\cdots,A_{1},\phi(S_{1})]$ $=$ $\mathbb{E}[R_{t}|A_{t},\phi(S_{t})],$
$\mathbb{P}(\phi(S_{t+1})=x|A_{t},\phi(S_{t}),A_{t-1},\phi(S_{t-1}),\cdots,A_{1},\phi(S_{1}))$ $=$ $\mathbb{P}(\phi(S_{t+1})=x|A_{t},\phi(S_{t})).$
The $\pi$-irrelevance condition in Definition 1 ensures that the Q-function $Q_{\phi}^{\pi}$, value function $V_{\phi}^{\pi}$ and IS ratios $\rho_{\phi}^{\pi}$, $w_{\phi}^{\pi}$ are well-defined on the abstract space. According to Definition 2, it is immediate to see that if $\phi$ is an MSA, then the transformed data tuples $(\phi(S),A,R,\phi(S^{\prime}))$ retain the MDP structure (we will discuss how to construct such MSAs in Section 4.2). This together with the $\pi$-irrelevance condition guarantees the identifiability of the nuisance functions $Q_{\phi}^{\pi}$, $V_{\phi}^{\pi}$ and $w_{\phi}^{\pi}$, meaning that they can be consistently estimated from the offline dataset. Below, we introduce a set of conditions to analyze the MSE of post-abstraction OPE estimators derived from these methods.
Assumption 1 (Boundedness)
The absolute value of all immediate rewards are uniformly bounded by a positive constant $R_{\mathrm{max}}<\infty$.
Assumption 2 (Coverage)
Both $\rho^{\pi}$ and $w^{\pi}$ are finite, i.e., $\rho^{\pi}(a,s),w^{\pi}(a,s)<\infty$ for any $a$ and $s$.
Assumption 3 (Stationary)
The state-action-reward triplets $(S_{t},A_{t},R_{t})_{t\geq 1}$ from the ground MDP are stationary over time.
Assumption 4 (Nuisance function estimators)
[…]
We make a few remarks: (i) Assumptions 1 and 2 are widely adopted in the RL and OPE literature (see, e.g., Chen and Jiang, 2019; Fan et al., 2020; Uehara et al., 2020; Yin and Wang, 2020). (ii) Assumption 3 is introduced for convenience to ensure that the denominator of the MIS ratio (see (2)) is well-defined. Notably, we do not require the state-action-reward-next-state tuples to be i.i.d., as in Munos and Szepesvári (2008); Sutton et al. (2008); Chen and Jiang (2019); Fan et al. (2020), since the independence assumption is clearly violated in MDPs. Similarly, we do not impose any restrictive ergodicity or mixing conditions on the ground MDP, as in Bhandari et al. (2018); Zou et al. (2019); Luckett et al. (2019); Shi et al. (2021); Chen and Qi (2022). (iii) The VC-class type condition in Assumption 4 is commonly used in machine learning and statistics (see, e.g., Van Der Vaart et al., 1996; Shalev-Shwartz and Ben-David, 2014) to measure the complexity of a model class. The boundedness conditions on the Q-function and ratio function classes are also mild. Given Assumptions (1) and (2), it follows immediately that the oracle Q-function and MIS ratio are bounded by $R_{\mathrm{max}}/(1-\gamma)$ and $O(1)$, respectively. It is thus reasonable to assume these function classes are bounded by these constants as well.
Theorem 1 (Bias-variance decomposition)
Assume Assumptions 14 hold. Then the followings hold for any $\pi$-irrelevant MSA $\phi$:
  • The MSE of DRL (Method 4) applied to the abstract MDP is given by
    $\begin{split}&\frac{1}{N}\mathbb{V}\left[V_{\phi}^{\pi}\left(\phi(S)\right)\right]+\frac{1}{NT(1-\gamma)^{2}}\mathbb{V}\Big{[}w_{\phi}^{\pi}(A,\phi(S))\big{[}R+\gamma V_{\phi}^{\pi}\left(\phi(S^{\prime})\right)-Q_{\phi}^{\pi}(A,\phi(S))\big{]}\Big{]}\\ &+O\Big{(}\frac{R_{\max}^{2}\max(\kappa_{q}(\phi),\kappa_{w}(\phi),N^{-1/2})\log(N)}{(1-\gamma)^{2}N}+\frac{R_{\max}^{2}\max(\kappa_{q}(\phi)\kappa_{w}(\phi),\sqrt{N}\kappa_{q}^{2}(\phi)\kappa_{w}^{2}(\phi))}{\sqrt{N}(1-\gamma)^{2}}\Big{)}.\end{split}$ (3)
    Here, $\mathbb{V}[\bullet]$ denotes the variance of a random variable, and we recall that $N$ denotes the number of trajectories, whereas $T$ denotes the number of state-action-reward-next-state tuples per trajectory.
  • When both the Q-function and MIS ratio are estimated via linear function approximation (refer to Appendix D.2 for the detailed estimation procedure), the MSEs of Q-function-based estimator (Method 1) and MIS (Method 3) applied to the abstract MDP equal to (3) as well.
  • When $\rho_{\phi}^{\pi}$ is known, the MSE of SIS (Method 2) applied to the abstract MDP is upper bounded by
    $\frac{4R_{\max}^{2}}{NT^{2}(1-\gamma)^{2}}\sum_{t=1}^{T}\gamma^{2t-2}\mathbb{E}(\rho_{\phi,1:t}^{\pi})^{2}+\frac{\gamma^{2T}R_{\max}^{2}}{(1-\gamma)^{2}}.$ (4)
At first glance, Theorem 1 may appear complex, but both the theorem and its proof have several important implications, as we elaborate below:
  1. 1.
    Equation (3) provides a bias-variance decomposition for the MSE of DRL. Specifically, in Equation (3), the first line represents the asymptotic variance of DRL, which is of order $O(N^{-1})$, while the second line upper bounds their squared finite-sample bias arising from the estimation of the Q-function and MIS ratio. Notably, when the estimation errors satisfy $\kappa_{q}(\phi)=O(N^{-\ell_{1}})$ and $\kappa_{w}(\phi)=O(N^{-\ell_{2}})$ for $\ell_{1},\ell_{2}>0$ and $\ell_{1}+\ell_{2}>1/2$ – conditions commonly imposed on doubly robust estimators for valid uncertainty quantification (see, e.g., Chernozhukov et al., 2018; Farrell et al., 2018; Kallus and Uehara, 2022)the bias term decays to zero at a much faster rate than the asymptotic variance term as $N$ increases to infinity.
  2. 2.
    When employing linear function approximation to model the nuisance functions, the resulting Q-function-based and MIS estimators become essentially equivalent to a DRL estimator (see Appendix D.2 for a formal proof). As a result, the same bias-variance decomposition in Equation (3) applies to the Q-function-based and MIS estimators as well.
  3. 3.
    Equation (4) provides a bias-variance decomposition for the MSE of SIS. Specifically, the two terms upper bound the variance and squared bias of SIS, respectively. Notably, the variance upper bound is proportional to the second moment of the product of the IS ratios, which reflects the curse of horizon. The bias term arises from truncating the data at time $T$, while the target policy to be evaluated involves an infinite sum. Similar to the other three estimators, the squared bias is typically much smaller than the variance.
Building on Theorem 1, we next investigate the impact of different abstractions on the bias and variance of the resulting OPE estimators. As shown in Equation (3), the bias term depends on the estimation errors of the Q-function and MIS ratio. To simplify the analysis, we use lookup tables to parameterize the Q-function and MIS ratio. For any abstraction $\phi$, let $\textrm{Var}^{(1)}(\phi)$ and $\textrm{Var}^{(2)}(\phi)$ denote the variance terms in Equations (3) and (4), respectively. The following theorem summarizes our findings.
Theorem 2 (Bias, variance and abstraction: The role of $\phi$ in OPE)
Let $\phi$ denote an MSA from $\mathcal{S}$ to $\mathcal{X}$ and $\phi^{*}$ denote another MSA from $\mathcal{X}$ to certain $\mathcal{X}^{*}=\{\phi^{*}(x):x\in\mathcal{X}\}$. Then we have:
  • $\lim_{T\to\infty}\textrm{Var}^{(1)}(\phi)\geq\lim_{T\to\infty}\textrm{Var}^{(1)}(\phi^{*}\circ\phi)$ and the equality holds if and only if $V^{\pi}_{\phi}$ is $\phi^{*}$-irrelevant, i.e., for any $x_{1},x_{2}\in\mathcal{X}$, $V_{\phi}^{\pi}(x_{1})=V_{\phi}^{\pi}(x_{2})$ whenever $\phi^{*}(x_{1})=\phi^{*}(x_{2})$.
  • When lookup tables are employed to parameterize the Q-function and MIS ratio, under certain matrix invertibility condition detailed in Appendix D, we have
    $\kappa_{q}(\phi)=O\Big{(}\frac{{|\mathcal{X}||\mathcal{A}|}}{\sqrt{NT}}\Big{)}\,\,\hbox{and}\,\,\kappa_{w}(\phi)=O\Big{(}\frac{{|\mathcal{X}||\mathcal{A}|}}{\sqrt{N}}\Big{)}.$
  • $\textrm{Var}^{(2)}(\phi)\geq\textrm{Var}^{(2)}(\phi^{*}\circ\phi)$ and the equality holds if and only if $\rho_{\phi}^{\pi}$ is $\phi^{*}$-irrelevant.
  • The bias term in (4) is a constant function of $\phi$.
Given that $\phi$ and $\phi^{*}\circ\phi$ are two nested abstractions, we summarize the implications of Theorem 2 as follows:
  1. 1.
    Results in the first bullet point suggest that the asymptotic variances of the Q-function-based estimator, MIS, and DRL decay with more refined abstractions for sufficiently large $T$.
  2. 2.
    Results in the second bullet point imply that the Q-function and MIS estimation errors generally decrease with the size of the state space. Combined with Equation (3), this suggests that the finite-sample biases of the Q-function-based estimator, MIS, and DRL also decay with more refined abstractions.
  3. 3.
    Results in the third bullet point suggest that the asymptotic variance of SIS decays with more refined abstractions.
  4. 4.
    Results in the last bullet point indicate that the bias of SIS is independent of the abstraction.
In summary, for all the four OPE estimators, their MSEs generally decrease with more refined abstractions. However, it is important to note that all the aforementioned results apply only to MSAs. When the abstraction fails to preserve the Markov property, the resulting OPE estimator will suffer from a large bias. The aforementioned discussions motivate us to identify the finest MSA for minimizing the resulting OPE estimator’s MSE. However, direct identification of such an MSA can be challenging. Instead, we propose an iterative state abstraction procedure, which identifies a sequence of MSAs $\phi_{1}$, $\phi_{2}$, $\cdots$, $\phi_{K}$ to progressive reduce the state space. Based on our theories, we expect the MSE of the resulting OPE estimator decreases with the number of iterations, i.e.,
$\textrm{MSE}(\phi_{1})\geq\textrm{MSE}(\phi_{2}\circ\phi_{1})\geq\cdots\geq\textrm{MSE}(\phi_{K}\circ\cdots\circ\phi_{1}),$
where the inequalities are likely to hold when these abstractions are not identical mappings. Thus, given a sufficiently large $K$, the accuracy of the OPE estimator can be substantially improved. We detail our methodology in the next section.
4.2 Iterative state abstraction
Summary of our approach. We begin with an overview of the proposed iterative procedure. As outlined earlier, our goal is to construct a sequence of abstractions $\phi_{1},\phi_{2},\dots,\phi_{K}$ to progressively compress the state space. Starting from $\mathcal{X}_{0}=\mathcal{S}$, the original state space, each $\phi_{k}$ maps from $\mathcal{X}_{k-1}$ to $\mathcal{X}_{k}$. The final abstraction is obtained by composing the sequence as $\phi_{K}\circ\cdots\circ\phi_{2}\circ\phi_{1}$, yielding a deeply-abstracted state space.
Based on our theoretical framework, these abstractions must satisfy the following requirements:
  1. 1.
    Each $\phi_{k}$ must be a $\pi$-irrelevant MSA to guarantee the bias-variance decompositions in Equations (3) and (4) hold.
  2. 2.
    Any two adjacent abstractions must employ different abstraction methods; otherwise, their composition may result in the same mapping that fails to reduce the state space.
To simultaneously satisfy both requirements, we observe that there are two types of MSAs available:
  • (i)
    The first is based on the notion of model-irrelevance (see e.g., Li et al., 2006); see also, Definition  5 below.
  • (ii)
    The second is the abstraction developed by Allen et al. (2021).
This motivates us to alternate between these two abstractions at each iteration of the procedure (see Figure 1 for an illustration). For both types of MSAs, we make necessary adjustments to adapt them to our setup. For instance, we tailor them to be $\pi$-irrelevant by concatenating them with $\{\pi(a|\bullet):a\in\mathcal{A}\}$ and extend Allen et al. (2021)’s abstraction by relaxing their requirement for the behavior policy to be Markovian —- an assumption likely violated due to the iterative nature of our procedure. We emphasize that the abstractions employed in each iteration of our procedure are not entirely new. Our primary methodological contribution lies in the iterative use of these abstraction, which, guided by our theoretical framework, can significantly enhance the performance of OPE. We next detail the two different types of MSAs used in our iterative procedure.
Figure 1: Illustrations of the iterative procedure.
Type-I MSA. We present model-irrelevance, a widely studied condition in the state abstraction literature, designed for learning the optimal policy $\pi^{*}$. To better understand this condition, we begin by reviewing some basic definitions of state abstraction for policy learning; see also Jiang (2018).
@@ 4.2 Iterative state abstraction
Definition 4 ($Q^{\pi^{*}}$-irrelevance)
$\phi$ is $Q^{\pi^{*}}$-irrelevant if for any $s^{(1)}$, $s^{(2)}\in\mathcal{S}$ whenever $\phi(s^{(1)})=\phi(s^{(2)})$, the optimal Q-function satisfies $Q^{\pi^{*}}(a,s^{(1)})=Q^{\pi^{*}}(a,s^{(2)})$ for any $a\in\mathcal{A}$.
Definitions 3 and 4 are easy to understand, requiring the optimal policy/Q-function to depend on a state $s$ only through its abstraction $\phi(s)$. However, the resulting $\phi$ may not necessarily preserve the Markov property. This leads to the following model-irrelevance, which preserves the MDP structure while ensuring $\pi^{*}$- and $Q^{*}$-irrelevance (Li et al., 2006).
@@
The first condition in (5) corresponds to “reward-irrelevance” whereas the second condition represents “transition-irrelevance”. Notice that the term $\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathcal{T}(s^{\prime}|a,s)$ – appearing in the second equation of (5) – represents the probability of transitioning to $\phi(S^{\prime})=x^{\prime}$ in the abstract state space. Thus, the second condition essentially requires the abstract next state $\phi(S^{\prime})$ to be conditionally independent of $S$ given $A$ and $\phi(S)$. Assuming $S$ can be decomposed into the union of relevant features and irrelevant features. The condition implies that the evolution of those relevant features depends solely on themselves, independent of those irrelevant features. This ensures that $\phi$ is an MSA, as shown in the following lemma.
Lemma 1
The abstraction derived from model-irrelevance (5) is an MSA.
To compute such a type-I MSA, we first adapt existing algorithms (Ha and Schmidhuber, 2018; François-Lavet et al., 2019; Gelada et al., 2019) to train a model-irrelevant abstraction $\phi$, parameterized via deep neural networks. We next combine $\phi(s)$ with $\{\pi(a|s):a\in\mathcal{A}\}$ to ensure $\pi$-irrelevance. Refer to Appendix B.1 for the detailed procedures.
@@
Type-II MSA. To illustrate the second type of MSA, we begin by discussing two perspectives of the data generated the MDP framework (see Figure 2 for graphical illustrations of both perspectives):
@@
    The first perspective is the traditional (forward) MDP model with all state-action-reward triplets sequenced by time index. This yields the model-based irrelevance condition defined in Definition 5.
@@
    The second perspective offers a backward view by reversing the time order. Specifically, due to the symmetric nature of the Markov assumption — implying that if the future is independent of the past given the present, the past must also be independent of the future given the present — the reversed state-action pairs also maintain the Markov property. Leveraging this property, we define another backward MDP model below, which motivates the second type of MSA.